# Interpreting Shared Circuits for Ordered Sequence Prediction in a Large Language Model

### v3

@@ Abstract
While transformer models exhibit strong capabilities on linguistic tasks, their complex architectures make them difficult to interpret. Recent work has aimed to reverse engineer transformer models into human-readable representations called circuits that implement algorithmic functions. We extend this research by analyzing and comparing circuits for similar sequence continuation tasks, which include increasing sequences of Arabic numerals, number words, and months. By applying circuit interpretability analysis, we identify a key sub-circuit in GPT-2 responsible for detecting sequence members and for predicting the next member in a sequence. Our analysis reveals that semantically related sequences rely on shared circuit subgraphs with analogous roles. Overall, documenting shared computational structures enables better model behavior predictions, identification of errors, and safer editing procedures. This mechanistic understanding of transformers is a critical step towards building more robust, aligned, and interpretable language models.
@@ 1 Introduction
Recent work in interpretability has uncovered transformer circuits that implement simple linguistic tasks, such as identifying indirect objects in sentences (Wang et al., 2022). However, only a few studies have focused on the existence of shared circuits (Merullo et al., 2023), in which circuits utilize the same sub-circuits for similar tasks. Identifying shared circuits assists in aligning AI via methods such as model editing (Meng et al., 2023), which precisely targets problematic areas for more efficient re-alignment without erroneously altering healthy components. Documenting the existence of shared circuits enables safer, more predictable model editing with fewer risks, as editing a circuit may affect another if they share sub-circuits (Hoelscher-Obermaier et al., 2023). Therefore, interpretability enables safer alignment by understanding adverse effect prevention.
@@
While models use the same components for different tasks, such as when there are far more tasks/features than neurons (Elhage et al., 2022), our focus is on locating components which are shared due to similar, re-usable functionality, and not for vastly different functionalities. Our work tackles the hypothesis that LLMs may re-use circuits across analogous tasks that share common abstractions. For instance, similar sequence continuation tasks, such as number words ("one two three") and months ("Jan Feb Mar"), can be analogously mapped to one another via the natural number abstraction (for example one and Jan are mapped to 1). As these tasks share a common abstraction, LLMs may have learned to efficiently re-use components that utilize shared patterns. Understanding how LLMs re-use components based on commonalities can shed light on how they represent and associate semantic concepts with one another (Gurnee and Tegmark, 2023). Not only would this enhance understanding of how LLMs actually perceive information, but it may have potential applications in transfer learning (Zhuang et al., 2020).
@@
Thus, in this paper, we demonstrate the existence of shared circuits for similar sequence continuation tasks, as the similarity across these tasks is clear, allowing us to cleanly pinpoint functionality. Our key finding is that there exist shared sub-circuits between similar tasks in GPT-2 (Radford et al., 2019), where the shared components have the same functionality across tasks. As shown in Figure 1, the circuit for continuing a sequence of numerals shares a sub-circuit with the circuit for continuing a sequence of number words, which generally handles sequence continuation functionality.
The main contributions of this work are: (1) The discovery of shared circuits for sequence continuation tasks, (2) Finding that similar tasks utilize sub-circuits with the same functionality, and (3) A simple iterative approach to find circuits at a coarse granularity level. This advances our understanding of the mechanisms of how transformer models generalize concepts by re-using components.
Figure 1: A simplified circuit showing important components for the Numerals (red) and Number Words (blue) tasks merged into one diagram. The purple portions denote a shared, entangled sub-circuit across both tasks.
@@ 2 Background and Related Work
Transformer Models. We analyze LLM transformer-based models that take in an input sequences $(\text{{x}}_{1},\ldots,\text{{x}}_{L})$ of length $L$ Tokens are mapped to d-dimensional embeddings (Vaswani et al., 2017). An embedding is a function $\Psi:\mathcal{D}\rightarrow\mathbb{R}^{d}$, mapping token elements from a vocabulary domain $\mathcal{V}$ to vectors in $\mathbb{R}^{d}$.
Attention Head. A transformer model consists of blocks of attention heads, which each consists of two matrices: the QK matrix that outputs the attention pattern $A_{i,j}$, and the OV matrix that outputs to the residual stream. The output of an attention layer is the sum of attention heads $h_{i,j}$. We use the notation L.H for attention heads, where L is a layer index and H is a head index in layer L.
An attention head is parameterized by key, query and value matrices $W_{k},W_{q},W_{v}\in\mathbb{R}^{d\times d}$. We denote the key, query and value outputs as $k_{i}=W_{k}\mathbf{x}_{i}$, $q_{i}=W_{q}\mathbf{x}_{i}$, $v_{i}=W_{v}\mathbf{x}_{i}$, and denote $K_{i}=[k_{1},\ldots,k_{i}]\in\mathbb{R}^{d\times i}$ and $V_{i}=[v_{1},\ldots,v_{i}]\in\mathbb{R}^{d\times i}$. The output of the head at token $i$ is denoted by $\mathbf{o}_{i}\in\mathbb{R}^{d}$, where $\mathbf{o}_{i}=V_{i}\cdot\text{softmax}(K_{i}\cdot q_{i})$.
@@
Transformer-Block. An attention-block combines information from multiple attention heads, with each head attending to different parts of the input sequence. In a transformer model, $l$ attention heads are applied in parallel, forming an attention-block. Subsequently, a transformer-block processes the output of the attention-block, typically followed by an MLP. The MLP operates on the concatenated output of the $l$ attention heads, allowing the model to capture complex patterns in the data.
@@
Circuit Discovery. To analyze computations within models, a recent approach has been to find circuits, which are subgraphs of neural networks that represent algorithmic tasks (Elhage et al., 2021). In transformer circuits, evidence has shown that in general, MLPs associate input information with features Geva et al. (2020), while attention heads move information (Olsson et al., 2022).
@@
Interpretability of Sequential Tasks. In concurrent work, successor heads were found by Gould et al. (2023) to increment tokens with a natural ordering. However, the authors do not look for circuits for sequences with more than one member, and only focus on components; in this paper, we also study shared interactions between components. Hanna et al. (2023) found circuits for "greater-than" sequence tasks (for example completing the sentence, “The war lasted from the year 1732 to the year 17”, with any valid two-digit end years > 32). Greater-than tasks allow any year greater than a value to be valid, which differs from our sequence completion tasks that only have one valid answer. The authors noted that "similar tasks had similar… circuits", but compared across number tasks, and not across non-number tasks such as months. In our work, we study tasks that are more dissimilar.
@@ 3.2 Functionality Discovery Methods
Attention Pattern Analysis. We analyze the $\mathbf{QK}$ matrix of attention heads to track information movement from keys to queries. We take the mean of datasets samples to calculate the attention scores of the heatmaps in §5.2 and §H, and display only one sample on the axes for demonstration.
Component Output Scores. We analyze head outputs by examining the values written to the residual stream via the heads’ output matrices ($\mathbf{OV}$), allowing us to see what information is being passed by each head along in the circuit. These values are measured by component output scores; we utilize a "next sequence" score that measures how well a head, given sequence token $I$, outputs token $I+1$. The details of this method are described in Appendix G.
@@
4 Discovering Circuit Connectivity
In this section, we describe the experimental setup for our ablation experiments. We observe that there are multiple circuits, with slight variations between them, that have similar performances for the same task. However, we find that important heads are often found in most circuits, regardless of the method, metric or dataset choices. Thus, we focus more on the "big picture" comparison of scores and on the most important heads, and less on the exact variations between scores or the less important heads. We ran experiments on a NVIDIA A100 GPU.
Model. We test on GPT-2 Small (117M parameters), which has 144 heads and 12 MLPs.
@@ 4 Discovering Circuit Connectivity
Datasets. We run a generated prompts dataset of length 4 sequences (eg. 1 2 3 4). We found that the model could continue Numerals sequences even past 1000. However, our focus in this paper is not on finding circuits only for Numeral sequences, but on prompt types that share a common abstraction. Thus, to better compare numbers to months, we use sequences ranging from 1 to 12.
For each task, we generate samples by placing our sequence members among non-sequence tokens. For instance, one sample may be ’Kyle was born in February. Anthony was born in March. Grant was born in April. Madison was born in’. Placing sequence members amongst non-sequence tokens allows us to evaluate the circuit representation of the shared sub-task of how the model selects sequence members from non-sequence members. We generate a total of 1536 samples per task; thus, there are 4608 total samples. More discussion about datasets is found in Appendix B.

### v4

# Towards Interpretable Sequence Continuation: Analyzing Shared Circuits in Large Language Models

@@ Abstract
While transformer models exhibit strong capabilities on linguistic tasks, their complex architectures make them difficult to interpret. Recent work has aimed to reverse engineer transformer models into human-readable representations called circuits that implement algorithmic functions. We extend this research by analyzing and comparing circuits for similar sequence continuation tasks, which include increasing sequences of Arabic numerals, number words, and months. By applying circuit interpretability analysis, we identify a key sub-circuit in both GPT-2 Small and Llama-2-7B responsible for detecting sequence members and for predicting the next member in a sequence. Our analysis reveals that semantically related sequences rely on shared circuit subgraphs with analogous roles. Additionally, we show that this sub-circuit has effects on various math-related prompts, such as on intervaled circuits, Spanish number word and months continuation, and natural language word problems. Overall, documenting shared computational structures enables better model behavior predictions, identification of errors, and safer editing procedures. This mechanistic understanding of transformers is a critical step towards building more robust, aligned, and interpretable language models.
**footnotetext: Equal contribution
@@ 1 Introduction
To analyze computations within models, a recent approach has been to find circuits, which are subgraphs of neural networks that represent algorithmic tasks (Elhage et al., 2021). Recent work in interpretability has uncovered transformer circuits that implement simple linguistic tasks, such as identifying indirect objects in sentences (Wang et al., 2022). However, only a few studies have focused on the existence of shared circuits (Merullo et al., 2023), in which circuits utilize the same sub-circuits for similar tasks. Identifying shared circuits assists in aligning AI via methods such as model editing (Meng et al., 2023), which precisely targets problematic areas for more efficient re-alignment without erroneously altering healthy components. Documenting the existence of shared circuits enables safer, more predictable model editing with fewer risks, as editing a circuit may affect another if they share sub-circuits (Hoelscher-Obermaier et al., 2023). Therefore, interpretability enables safer alignment by understanding adverse effect prevention.
@@
While models use the same components for different tasks, such as when there are far more tasks/features than neurons (Elhage et al., 2022), our focus is on locating components which are shared due to similar, re-usable functionality, and not for vastly different functionalities. Our work tackles the hypothesis that LLMs may re-use circuits across analogous tasks that share common abstractions. For instance, similar sequence continuation tasks, such as number words ("one two three") and months ("Jan Feb Mar"), can be analogously mapped to one another via the natural number abstraction (for example, one and Jan are mapped to 1). As these tasks share a common abstraction, LLMs may have learned to efficiently re-use components that utilize shared patterns. Understanding how LLMs re-use components based on commonalities can shed light on how they represent and associate semantic concepts with one another (Gurnee and Tegmark, 2023). Not only would this enhance understanding of how LLMs actually perceive information, but it may have potential applications in transfer learning (Zhuang et al., 2020).
@@
Thus, in this paper, we demonstrate the existence of shared circuits for similar sequence continuation tasks, as the similarity across these tasks is clear, allowing us to cleanly pinpoint functionality. Our key finding is that there exist shared sub-circuits between similar tasks in GPT-2 Small (Radford et al., 2019) and Llama-2-7B (Touvron et al., 2023), where the shared components have the same functionality across tasks.
The main contributions of this work are:
  1. 1)
    The discovery of shared circuits for similar sequence continuation tasks.
  2. 2)
    The finding of similar sub-circuits across models with analogous functionality.
  3. 3)
    Showing that these circuits affect natural language math-related prompts.
These contributions improve the understanding of how transformer models generalize concepts by re-using components for analogous tasks.
Figure 1: The important components of a shared, entangled sub-circuit for the Numerals, Number Words, and Months tasks in GPT-2 Small. The functional roles of the components are labeled below them. Resid_post denotes the residual stream state right before the linear unembedding to logits. Full circuits are shown in Appendix H.
@@ 2 Background and Related Work
Transformer Models. A transformer model learns the importance of different parts of the data in relation to each other via query-key attention mechanisms. We analyze LLM transformer-based models that take in text inputs as token sequences $(x_{1},\ldots,x_{L})$ of length $L$. Tokens are mapped to $d$-dimensional embeddings by a function $\Psi:\mathcal{D}\rightarrow\mathbb{R}^{d}$ (Vaswani et al., 2017).
Attention Head. A transformer consists of blocks of attention heads, each consisting of two matrices: the query-key QK matrix that outputs the attention pattern $A_{i,j}$, and the output-value OV matrix that outputs to the residual stream. An attention layer output is the sum of attention heads $h_{i,j}$. We use the notation L.H for attention heads, where L is a layer index and H is a head index in layer L.
@@
Transformer Circuit Discovery. In transformer circuits, evidence has shown that in general, MLPs associate input information with features Geva et al. (2020), while attention heads move information (Olsson et al., 2022).
@@
Interpretability of Sequential Tasks. In concurrent work, successor heads were found by Gould et al. (2023) to increment tokens with a natural ordering. However, Gould et al. (2023) only analyze the output of the successor head, rather than looking at the output of a model, and does not look for circuits for sequences with more than one member. As shown in our paper, a successor head is not enough to evaluate a sequence. Additionally, we study shared interactions between components. Thus, we build on Gould et al. (2023) by showing how successor heads interact with the model as a whole. Hanna et al. (2023) found circuits for "greater-than" sequence tasks (for example completing the sentence, “The war lasted from the year 1732 to the year 17”, with any valid two-digit end years > 32). Greater-than tasks allow any year greater than a value to be valid, which differs from our sequence completion tasks that only have one valid answer. The authors noted that "similar tasks had similar… circuits", but compared across number tasks, and not across non-number tasks such as months. In our work, we study tasks that are more dissimilar. Related to sequence continuation, Stolfo et al. (2023) and (Quirke and Barez, 2023) discover circuits for arithmetic computation.
@@ 3.2 Functionality Discovery Methods
Attention Pattern Analysis. We analyze the $\mathbf{QK}$ matrix of attention heads to track information movement from keys to queries. We take the mean of datasets samples to calculate the attention scores of the heatmaps in §5.2 and §L, and display only one sample on the axes for demonstration.
Component Output Scores. We analyze head outputs by examining the values written to the residual stream via the heads’ output matrices ($\mathbf{OV}$), allowing us to see what information is being passed by each head along in the circuit. These values are measured by component output scores; we utilize a "next sequence" score that measures how well a head, given sequence token $I$, outputs token $I+1$. The details of this method are described in Appendix K.
@@
4 Circuit Connectivity in GPT-2 Small
4.1 Experimental Setup
In this section, we describe the experimental setup for our ablation experiments. We ran experiments on a NVIDIA A100 GPU. More setup details can be found in Appendix F.
Models. We test on GPT-2 Small (117M parameters), which has 144 heads and 12 MLPs
@@ 4.1 Experimental Setup
Datasets. For GPT-2 Small, we run a generated prompts dataset of length 4 sequences (eg. 1 2 3 4). Our focus in this paper is not on finding circuits only for Numeral sequences, but on prompt types that share a common abstraction. Thus, to better compare numbers to months, we use sequences ranging from 1 to 12.
For each task, we generate samples by placing our sequence members among non-sequence tokens. For instance, one sample may be ’Kyle was born in February… Grant was born in April. Madison was born in’. Placing sequence members amongst non-sequence tokens allows us to evaluate the circuit representation of the shared sub-task of how the model selects sequence members from non-sequence members. We generate a total of 1536 samples per task, with a total of 4608 samples.